### v1

@@ 3 Methods
Imaging from a single spacecraft cannot directly derive the 3D structure of a CME since it only captures a 2D projection in the plane of the image. Several techniques for CME 3D reconstructions, especially for their extended fronts, were developed during the STEREO era, which offered at least two different vantage points from where CMEs could be observed (see, e.g., Lugaz et al., 2009; Temmer et al., 2009; Byrne et al., 2010; Mierla et al., 2010; Rollett et al., 2016). The most widely used method is the so-called Graduated Cylindrical Shell (GCS) model (Thernisien et al., 2009; Thernisien, 2011). The parameters used to define the GCS model are the longitude and latitude at the apex of the CME, its height, the rate of expansion $\kappa$, the half-angle $\alpha$, which represents the angle between the main axes and half of one leg, and the tilt with respect to the ecliptic plane. We applied GCS on stereoscopic image data from STA and SoHO, both located at about 1 AU (separation angle $\approx$ ${}^{\circ}$), and PSP. The software package PyThea (Kouloumvakos et al., 2022) was used to perform the GCS reconstruction on the HI-1 data, while the package developed by Von Forstner (2021) was used for the reconstruction on COR2, C3, and WISPR data.
As PSP is moving quite fast (around 160 km/s near the E14 perihelion), the single spacecraft image data can be used to derive the 3D morphology and kinematics. The method is described in Liewer et al. (2020) and details of its application in this study are given in the Appendix A. It assumes that the structures move radially outwards at a constant velocity. Another requirement is that the spacecraft angular position varies sufficiently, for which we take a minimum of ${}^{\circ}$. Using Equations 3 and 4 we derive the velocity, position, latitude and longitude of a tracked small-scale structure. To solve the equations, we follow the approaches of both Liewer et al. (2020) and Braga & Vourlidas (2021). Uncertainties are estimated by tracking each identified feature at least three times.
@@ 4.1 Small-scale structure morphology
Figure 2: Selected LW-processed snapshots of the WISPR-I WT data set obtained on December 8, 2022 between 15 UT and 20 UT showing the tracked features (labeled 0 through 4). The features tracked belong to structures resembling a J-shape (0) or C-shape (1, 2) as well as being fishbone-like revealing a inclined thread-like pattern (3, 4). The x-axis shows the helioprojective longitude, the y-axis shows the helioprojective latitude. The animation corresponding to this figure is available online in the file movie1.mov.
@@
Figure 3: WISPR-I images from December 8 20:30 UT to December 9 18:30 UT and tracked features (0–9). The information from these images are complementary to those from Figure 2.
@@
Figure 4: The features 6–9 (blob-like) propagate in the wake of the CME. The information from these images are complementary to those from Figure 2 and 3.
@@
In Figures 24, we display snapshots of the selected small-scale structures, which we observe as the CME evolves within the FoV of WISPR-I and WISPR-O. All the tracked small-scale structures have been carefully selected considering a significant variation of the HCI longitude of PSP, together with a sufficient visibility/contrast that they need to have over time (see Appendix A). In total we tracked 10 different features over a time of about 32 hours (labeled 0 through 9, in chronological order in which they entered the WISPR-I FoV). In particular, Figure 2 shows the tracked features 0–4 of the CME in a section of the FoV of WISPR-I, when the instrument is in the WT mode (see Section 2). The tracked features clearly outline curved geometries, which might be related to twisted magnetic field structures. A J-shape structure, that resembles half of a U-shape, is well visible; therefore, we perform a track at its upper edge (feature 0), which stays longer in both FoVs of WISPR-I and WISPR-O with respect to the lower edge. Multiple C-shaped structures are also observed, with track features 1 and 2 located on the “C” upper and lower edges of “C ”. Also present is a well-defined thread-like structure (features 3 and 4), typical for filaments.
@@
Figure 3 shows the small-scale structures captured by WISPR-I in synoptic mode. Interesting to note is another C-shape feature, namely feature 5. As the synoptic mode covers a wider FoV, the left panel of Figure 3 reveals the total of features 0–5. Figure 4 shows a later time sequence on 9 December 2022 and the aftermath of the CME. Here a series of blobs (features 6, 7, 8, 9) is visible propagating in the wake of the CME. As time progresses, we notice here how the C-shaped structure at the front expands over time and changes its shape (compare with Figure 2). Because the CME is directed southward, many features tracked in WISPR-I do not reach WISPR-O. In addition, the features become fainter as they move to the outer instrument. As a consequence, only features 0 and 2 can be clearly tracked in the FoV of WISPR-O, as shown in Figure 5. All other features have been tracked using WISPR-I.
@@
Figure 5: WISPR-O images on December 8 23:39 UT and December 9 01:39 UT with tracked features 0 and 2.
@@ 4.2 Small-scale structure 3D single-spacecraft reconstruction
Over the duration of the event from 15:30 UT on December 8 to 23:30 UT on December 9, PSP longitude varies by about ${}^{\circ}$ (HCI; cf.  left panel of Figure 1). Each of the ten features was tracked multiple times with at least 3 repetitions and the methods by Braga & Vourlidas (2021) and by (Liewer et al., 2020) were applied to derive their 3D trajectories. Appendix A covers all related plots for deriving the results using the two different approaches. The results are summarized for both methods in Table 1, which gives for each feature the HCI longitude, latitude as well as initial speed and distance. Additionally, we report the angular distance, $\Delta\phi_{1}$, PSP has traveled in longitude (HCI) during the tracking of that specific feature, the angular distance $\phi_{2}-\phi_{1}$ between the feature and PSP, and finally the morphological group.
Within the uncertainties, we obtain for the tracked features comparable results for all the parameters derived from the two different methods. All the results shown are in HCI coordinate system. In general, all features propagate within a longitudinal range from ${}^{\circ}$ to ${}^{\circ}$ and a latitudinal range from ${}^{\circ}$ to ${}^{\circ}$. Inspecting the results in Table 1, we find that features 0, 1 and 2 propagate along ${}^{\circ}$ to ${}^{\circ}$ in longitude and between ${}^{\circ}$ and ${}^{\circ}$ in latitude with average speeds of $\sim$480 km/s. Feature 5 propagates in a similar direction as features 0–2 (longitude from ${}^{\circ}$ to ${}^{\circ}$ and latitude from ${}^{\circ}$ to ${}^{\circ}$), though likely having a lower speed ($\sim$450 km/s). Features 3 and 4 propagate along a clearly different direction (longitude between ${}^{\circ}$ and ${}^{\circ}$, latitude from ${}^{\circ}$ to ${}^{\circ}$) with speeds between 300–400 km/s. The blob-like features 6–9 propagate at similar speeds ($\sim$300-400 km/s) and latitudes from ${}^{\circ}$ to ${}^{\circ}$, but are distributed over a large longitudinal range between ${}^{\circ}$ and ${}^{\circ}$.
By their similarity in morphology and dynamics, the tracked features can be divided into three groups: group 1 (features: 0, 1, 2, and 5) most likely refers to twisted magnetic field structures with C- and J-shapes, group 2 (features: 3 and 4) to thread-like patterns and group 3 (features, 6, 7, 8, and 9) to blobs propagating for several hours in the rear part of the CME. On average we see a decreasing trend in speed from the fastest group 1 to the slowest group 3. While each feature of group 1 and 2 shows a clustered behavior, i.e., reveal roughly similar speed, longitude and latitude among each group, the blobs span a wide range of longitudes ($\sim$ ${}^{\circ}$).
@@
ID
$t_{0}$
$\Delta\phi_{1}$ (deg)
$\phi_{2}$ (deg)
$\phi_{2}$-$\phi_{1}$ (deg)
$\delta_{2}$ (deg)
$v_{0}$ (km/s)
$r_{0}$ (Rdirect-$\odot$)
group
0
2022-12-08T15:00:45.635
6.2
$22\pm 3$
@@
Table 1: Results in HCI obtained by applying the methods discussed in Section 3 to the features labeled in Figures 2 and 3. ID is the label of the feature, $t_{0}$ is the initial time at which each feature is tracked, $\Delta\phi_{1}$ is the variation in the longitude of the PSP spacecraft during the tracking of the specific feature, $\phi_{2}$ and $\delta_{2}$ is the derived longitude and latitude, respectively, while $v_{0}$ and $r_{0}$ are the initial velocity and position of the feature. The last column represents the group in which the different features belong, see Section 4.2. In blue we show the results obtained with the fitting method of Braga & Vourlidas (2021), while in red the one of Liewer et al. (2020) (see Appendix A for more details).
@@ 4.3 Global GCS 3D reconstruction
Comparison to the 3D reconstruction of the CME volume allows us to interpret in more detail where the tracked small-scale structures are located with respect to the global appearance of the CME. The free parameters of the GCS model, the apex height of the idealized flux rope, angular width, and aspect ratio, are fitted so as to match the global structure of the CME in stereoscopic images from C3 and COR2. Figure 6 shows the GCS reconstruction at $\sim$16:30 UT on December 8, 2022. The GCS fit indicates a propagation direction for the apex roughly along 52° in longitude and --42° in latitude (HCI coordinates) with uncertainties in the range of about 10° for latitude and longitude, respectively (for detailed estimates on the accuracy of the GCS method we refer to Verbeke et al., 2023). The GCS geometry is described by the results $\alpha$=22°, $\kappa$=0.26 rad and a tilt of ${}^{\circ}$. From this we calculate the angular width face-on ($\omega_{\rm FO}$), which describes the lateral extension of the CME, by $\omega_{\rm FO}$=2($\alpha$+$\delta$) with $\delta=\sin^{-1}\kappa$, together with the edge-on angular width $\omega_{EO}=2\delta$ (see Table 1 in Thernisien, 2011). According to this, the CME has a half-width $\omega_{\rm FO}$ of $\sim$37° and an angular width at the edge $\omega_{EO}$ of $\sim 30^{\circ}$. Therefore, the CME is spanning from about 15 to 89 in HCI longitude and from --12 to --72 in latitude.
@@
Next, we expand the GCS fit to the inner Heliosphere combining HI-1 and WISPR-I data [^2] We note that a $\sim$7 minute time delay between observations from 1 AU (STA) and from 0.11-0.16 AU (PSP) should be considered to account for the light travel time difference between PSP and the spacecraft at 1 AU.. For that we used the GCS results derived from stereoscopic image data combining C3 and COR2 at $\sim$ 16:30 UT, kept all other parameters constant, and simply changed the height over time to match the CME front as observed in the heliospheric image data pairs. Figure 7 shows the GCS model applied to images from HI-1 and WISPR-I, at $\sim$17:00 UT ($H$ = 24 $R_{\odot}$), $\sim$19:30 UT ($H$ = 36 ${}_{\odot}$), and $\sim$23:22 UT ($H$ = 50 $R_{\odot}$). With this we derive a 3D CME speed of about 650 km/s. We observed distinct differences in the appearance of the CME between WISPR and HI-1. The WISPR images contain more small-scale structures, but the broad outer envelope is missing. The latter is clearly visible in both COR2 and HI-1.
Figure 6: GCS reconstruction on the running difference coronagraph images of COR2 and C3 at $\sim$16:30 UT.
@@
Figure 7: GCS reconstruction on the heliospheric images of HI-1 and WISPR 1 at three instant of time.
@@
To visualize the location of the different features in 3D and relate the results in Table 1 with those of GCS, we calculate the position of each small-scale structure at 16:30 UT on December 9, 2022 ($t_{0}$ of feature 9). Using the derived 3D speed, we extrapolate the height of the GCS apex accordingly to 110 ${}_{\odot}$ for the same time. Figure 8 shows the resulting 3D distribution of the features with respect to the central axis of the GCS located at $\Delta\phi=0$ deg, $\Delta\delta=0$ deg and spanning 0–110 ${}_{\odot}$ in height. The tracked small-scale structures are shown in boxes whose color depends on the morphological group discussed in Section 4.2. Features 0–7 are all found to lie within the GCS reconstructed volume and can therefore be safely interpreted as being CME internal small-scale structures. Comparing the location of the different groups, as listed in Table 1, to the central axis of the GCS, we find that features 0, 1, 2 and 5 are located in the eastern leg of the CME, while features 3 and 4 are part of the western leg.
@@
Figure 8: Comparison of the positions of the local magnetic structures, summarized in Table 1, together with the global CME flux rope modeled by GCS. All panels show 3D plots in a $\Delta\phi$, $\Delta\delta$, $R$ plane at 16:30 UT on December 9, 2022. $\Delta\phi$ and $\Delta\delta$ are the relative variation in longitude and latitude with respect to the central axis of the GCS, while R is the height. W indicates the west, E the east, S the south and N the north. The eye represents the point of view of the observer. The colors of the boxes indicate the different families to which the features 0-9 belong: 1 (azure), 2 (yellow) and 3 (gray). The plot shows: (A) the frontal view of the GCS modeled flux-rope and the relative location of the features with respect to its central axis, (B) shows the extension of the flux rope from a top view, (C) shows it from a west-top view, (D) shows it from the western leg. An animation of this image is available in the online material of this article.

### v2

@@ 3 Methods
Imaging from a single spacecraft cannot directly derive the 3D structure of a CME since it only captures a 2D projection in the plane of the image. Several techniques for CME 3D reconstructions, especially for their extended fronts, were developed during the STEREO era, and offered at least two different vantage points from where CMEs could be observed (see, e.g., Lugaz et al., 2009; Temmer et al., 2009; Byrne et al., 2010; Mierla et al., 2010; Rollett et al., 2016). The most widely used method is the so-called graduated cylindrical shell (GCS) model (Thernisien et al., 2009; Thernisien, 2011). The parameters used to define the GCS model are the longitude and latitude at the apex of the CME; its height; the rate of expansion $\kappa$; the half angle $\alpha$, which represents the angle between the main axes and half of one leg; and the tilt with respect to the ecliptic plane. We applied the GCS model to stereoscopic image data from STA and SoHO, both located at about 1 AU (separation angle $\approx$ 14), and PSP. The software package PyThea (Kouloumvakos et al., 2022) was used to perform the GCS reconstruction on the HI-1 data, while the package developed by Von Forstner (2021) was used for the reconstruction on COR2, C3, and WISPR data.
As PSP moves quite fast (around 160 km/s near the E14 perihelion), the single spacecraft image data can be used to derive the 3D morphology and kinematics. The method is described in Liewer et al. (2020), and details of its application in this study are given in Appendix A. The method assumes that the structures move radially outward at a constant velocity. Another requirement is that the spacecraft angular position varies sufficiently, for which we take a minimum of 4. Using Eqs. (3)-(4), we derived the velocity, position, latitude, and longitude of a tracked small-scale structure. To solve the equations, we followed the approaches of both Liewer et al. (2020) and Braga & Vourlidas (2021). Uncertainties were estimated by tracking each identified feature at least three times.
@@ 4.1 Small-scale structure morphology
Figure 2: Selected LW-processed snapshots of the WISPR-I WT data set obtained on December 8, 2022, between 15 UT and 20 UT showing the tracked features (labeled 0 through 4). The tracked features belong to structures resembling J-shapes (0), C-shapes (1, 2), and fish-bone-like shapes revealing an inclined thread-like pattern (3, 4). The x-axis shows the helioprojective longitude; the y-axis shows the helioprojective latitude. The animation corresponding to this figure is available online in the file movie1.mov.
@@
Figure 3: WISPR-I images from December 8 20:30 UT to December 9 18:30 UT and tracked features (0–9). The information from these images is complementary to that of Fig 2.
@@
Figure 4: Features 6 to 9 (blob-like) propagate in the wake of the CME. The information from these images is complementary to that from Figs. 2-3.
@@
In Figs. 24, we display snapshots of the selected small-scale structures, which we observed as the CME evolved within the FoV of WISPR-I and WISPR-O. All the tracked small-scale structures have been carefully selected, considering a significant variation of the HCI longitude of PSP, together with the sufficient visibility/contrast that they need to have over time (see Appendix A). In total we tracked ten different features over a period of about 32 hours (labeled 0 through 9, in the chronological order in which they entered the WISPR-I FoV). In particular, Fig. 2 shows the tracked features 0–4 of the CME in a section of the FoV of WISPR-I, when the instrument was in the WT mode (see Sect.  2). The tracked features clearly outline curved geometries, which might be related to twisted magnetic field structures. A J-shaped structure, that resembles half of a U-shape, is clearly visible; therefore, we performed a track at its upper edge (feature 0), that stays longer in both FoVs of WISPR-I and WISPR-O with respect to the lower edge. Multiple C-shaped structures were also observed, with track features 1 and 2 located on the upper and lower edges of the ”C” shape. Also present is a well-defined thread-like structure (features 3 and 4), which is typical for filaments.
@@
Figure 3 shows the small-scale structures captured by WISPR-I in synoptic mode. An interesting item to note is another C-shaped feature, namely feature 5. As the synoptic mode covers a wider FoV, the left panel of Fig. 3 reveals the total of features 0–5. Figure 4 shows a later time sequence on December 9, 2022, and the aftermath of the CME. Here, a series of blobs (features 6, 7, 8, 9) is visible, propagating in the wake of the CME. As time progresses, we noticed here how the C-shaped structure at the front expands over time and changes its shape (compared to Fig. 2). Because the CME is directed southward, many features tracked in WISPR-I do not reach WISPR-O. In addition, the features become fainter as they move to the outer instrument. As a consequence, only features 0 and 2 can be clearly tracked in the FoV of WISPR-O, as shown in Fig. 5. All other features were tracked using WISPR-I.
@@
Figure 5: WISPR-O images from December 8 23:39 UT and December 9 01:39 UT with tracked features 0 and 2.
@@ 4.2 Small-scale structure 3D single-spacecraft reconstruction
Over the duration of the event from 15:30 UT on December 8 to 23:30 UT on December 9, PSP longitude varied by about $\sim$22 (HCI; cf.  left panel of Fig.  1). Each of the ten features was tracked multiple times with at least three repetitions and the methods by Braga & Vourlidas (2021) and by (Liewer et al., 2020) were applied to derive their 3D trajectories. Appendix A covers all related plots for deriving the results using the two different approaches. The results are summarized for both methods in Table 1, which gives the HCI longitude and latitude as well as the initial speed and distance from the Sun for each feature. Additionally, we report the angular distance, $\Delta\phi_{1}$, PSP traveled in longitude (HCI) during the tracking of that specific feature; the angular distance, $\phi_{2}-\phi_{1}$, between the feature and PSP; and finally the morphological group.
Within the uncertainties, we obtained for the tracked features comparable results for all the parameters derived from the two different methods. All the results shown are in the HCI coordinate system. In general, all features propagate within a longitudinal range from 17 to 76 and a latitudinal range from --23 to --45. Inspecting the results in Table 1, we found that features 0, 1, and 2 propagate along 17 to 26 in longitude and between --23 and --30 in latitude, with average speeds of $\sim$480 km/s. Feature 5 propagates in a similar direction to features 0–2 (longitude from 31 to 38 and latitude from --36 to --37), though likely with a lower speed ($\sim$450 km/s). Features 3 and 4 propagate along a clearly different direction (longitude between 69 and 76, latitude from --43 to --45) with speeds between 300–400 km/s. The blob-like features 6–9 propagate at similar speeds ($\sim$300-400 km/s) and latitudes (from --31 to --39) but are distributed over a large longitudinal range between 24 and 66.
Based on their similarity in morphology and dynamics, the tracked features can be divided into three groups. Group 1 (features: 0, 1, 2, and 5) most likely refers to twisted magnetic field structures with C- and J-shapes. Group 2 (features: 3 and 4) refers to thread-like patterns, and group 3 (features, 6, 7, 8, and 9) refers to blobs propagating for several hours in the rear part of the CME. On average, we observed a decreasing trend in speed from the fastest group (group 1) to the slowest (group 3). While each feature of group 1 and 2 shows a clustered behavior, that is, they reveal a roughly similar speed, longitude, and latitude among each group, the blobs (group 3) span a wide range of longitudes ($\sim$ 40).
Table 1: Results in HCI obtained by applying the methods discussed in Sect. 3 to the features labeled in Figs. 2-3.
@@
ID $t_{0}$ $\Delta\phi_{1}$ (deg) $\phi_{2}$ (deg) $\phi_{2}$-$\phi_{1}$ (deg) $\delta_{2}$ (deg) $v_{0}$ (km/s) $r_{0}$ (Rdirect-$\odot$) group
0 2022-12-08T15:00:45.635 6.2 $22\pm 3$
@@
 [^2] The label of the feature, is indicated by its ID; $t_{0}$ is the initial time at which each feature was tracked; $\Delta\phi_{1}$ is the variation in the longitude of the PSP spacecraft during the tracking of the specific feature; $\phi_{2}$ and $\delta_{2}$ are the derived longitude and latitude, respectively; and $v_{0}$ and $r_{0}$ are the initial velocity and position of the feature. The last column represents the group in which the different features belong (see Sect. 4.2). In blue we show the results obtained with the fitting method of Braga & Vourlidas (2021), while in red we show the results from the fitting method of Liewer et al. (2020) (see Appendix A for more details).
@@ 4.3 Global GCS 3D reconstruction
Comparison to the 3D reconstruction of the CME volume allowed us to interpret in more detail where the tracked small-scale structures are located with respect to the global appearance of the CME. The free parameters of the GCS model (the apex height of the idealized flux rope, the angular width, and aspect ratio) were fitted so as to match the global structure of the CME in stereoscopic images from C3 and COR2. Figure 6 shows the GCS reconstruction at $\sim$16:30 UT on December 8, 2022. The GCS fit indicates a propagation direction for the apex roughly along 52° in longitude and --42° in latitude (HCI coordinates) with uncertainties in the range of about 10° for latitude and longitude, respectively (for detailed estimates on the accuracy of the GCS method we refer to Verbeke et al., 2023). The GCS geometry is described by the results $\alpha$=22°, $\kappa$=0.26 rad, and a tilt of --52. From this, we calculated the angular width face-on ($\omega_{\rm FO}$), which describes the lateral extension of the CME, by $\omega_{\rm FO}$=2($\alpha$+$\delta$) with $\delta=\sin^{-1}\kappa$, as well as the edge-on angular width $\omega_{EO}=2\delta$ (see Table 1 in Thernisien, 2011). According to the GCS modelled flux rope, the CME has a half width, $\omega_{\rm FO}$, of $\sim$37° and an angular width at the edge $\omega_{EO}$ of $\sim 30^{\circ}$. Therefore, the CME spans from about 15 to 89 in HCI longitude and from --12 to --72 in latitude.
@@
Next, we expand the GCS fit to the inner Heliosphere by combining HI-1 and WISPR-I data. [^3] We note that a $\sim$7 minute time delay between observations from 1 AU (STA) and from 0.11-0.16 AU (PSP) should be considered to account for the light travel time difference between PSP and the spacecraft at 1 AU. For that we used the GCS results derived from the stereoscopic image data combining C3 and COR2 at $\sim$ 16:30 UT, kept all other parameters constant, and simply changed the height over time to match the CME front as observed in the heliospheric image data pairs. Figure 7 shows the GCS model applied to images from HI-1 and WISPR-I at $\sim$17:00 UT ($H$ = 24 $R_{\odot}$), $\sim$19:30 UT ($H$ = 36 R), and $\sim$23:22 UT ($H$ = 50 $R_{\odot}$). With this approach, we derive a 3D CME speed of about 650 km/s. We observed distinct differences in the appearance of the CME between WISPR and HI-1. The WISPR images contain more small-scale structures, but the broad outer envelope is missing. The latter is clearly visible in both COR2 and HI-1.
Figure 6: GCS reconstruction of the running difference coronagraph images of COR2 and C3 at $\sim$16:30 UT.
@@
Figure 7: GCS reconstruction of the heliospheric images of HI-1 and WISPR 1 at three instants of time.
@@
To visualize the location of the different features in 3D and relate the results in Table 1 with those of GCS, we calculated the position of each small-scale structure at 16:30 UT on December 9, 2022 ($t_{0}$ of feature 9). Using the derived 3D speed, we extrapolated the height of the GCS apex accordingly to 110 R for the same time. Figure 8 shows the resulting 3D distribution of the features with respect to the central axis of the GCS located at $\Delta\phi=0$ deg, $\Delta\delta=0$ deg and spanning 0–110 R in height. The tracked small-scale structures are shown in boxes whose color depends on the morphological group discussed in Sect.  4.2. Features 0–7 were all found to lie within the GCS reconstructed volume and can therefore be safely interpreted as being CME internal small-scale structures. Comparing the location of the different groups, as listed in Table 1, to the central axis of the GCS, we found that features 0, 1, 2, and 5 are located in the eastern leg of the CME, while features 3 and 4 are part of the western leg.
@@
Figure 8: Comparison of the positions of the local magnetic structures summarized in Table 1 together with the global CME flux rope modeled by GCS. All panels show 3D plots in a $\Delta\phi$, $\Delta\delta$, $R$ plane at 16:30 UT on December 9, 2022. The relative variation in longitude and latitude with respect to the central axis of the GCS are represented by $\Delta\phi$ and $\Delta\delta$, while R is the height. The letter W indicates the west, E is east, S is south and N is north. The eye represents the point of view of the observer. The colors of the boxes indicate the different families to which the features 0 to 9 belong: 1 (azure), 2 (yellow), and 3 (gray). The plot shows the following: (A) the frontal view of the GCS modeled flux rope and the relative location of the features with respect to its central axis; (B) the extension of the flux rope from a top view; (C) same as plot B but from a west-top view, (D) same as plot B but from the western leg.